# hyperrefToken not allowed in a PDF string

### v1

@@ 1.1 Related work
In the presence of constraints that the final solution needs to satisfy, but which we can violate during exploration, Gelbart et al. (2014) propose to combine typical BO acquisition functions with the probability of satisfying the constraint. Instead, Gotovos et al. (2013) propose an uncertainty-based criterion that learns about the feasible region of parameters, while Perrone et al. (2019) modify the Max-value Entropy Search selection criterion (Wang and Jegelka, 2017) to include information about the constraint as well in the acquisition function. When we are not allowed to ever evaluate unsafe parameters, safe exploration is a necessary sub-routine of BO algorithms to learn about the safety of parameters. To safely explore, Schreiter et al. (2015) globally learn about the constraint by evaluating the most uncertain parameters. SafeOpt by Sui et al. (2015) extends this to joint exploration and optimization and makes it more efficient by explicitly restricting safe exploration to the boundary of the safe set. Sui et al. (2018) proposes StageOpt, which additionally separates the exploration and optimization phases. Both of these algorithms assume access to a Lipschitz constant to define parameters close to the boundary of the safe set, which is a difficult tuning parameter in practice. These methods have been extended to multiple constraints by Berkenkamp et al. (2016a), while Kirschner et al. (2019) scale them to higher dimensions with LineBO, which explores in low-dimensional sub-spaces. To improve computational costs, Duivenvoorden et al. (2017) suggest a continuous approximation to SafeOpt without providing exploration guarantees. Chan et al. (2023), on the other hand, take a more directed approach, by searching the most promising parameter also outside of the current safe set and then evaluating the most uncertain safe parameter in the neighborhood of the projection of such promising parameter on the boundary of the safe set. In the context of safe active learning, Li et al. (2022) consider the safe active learning problem for a multi-output Gaussian process, and choose the parameters to evaluate as those with the highest predictive entropy and high probability of being safe. All of these methods rely on function uncertainty to drive exploration, while we directly maximize the information gained about the safety of parameters.
@@ 2 Problem Statement and Preliminaries
In this section, we introduce the problem and the notation that we use throughout the paper. We are given an unknown and expensive function $f\mathrel{\mathop{:}}\mathcal{X}\rightarrow\mathbb{R}$ that we aim to optimize by sequentially evaluating it at parameters $\bm{x}_{n}\in\mathcal{X}$. However, we are not allowed to evaluate parameters that violate some notion of safety. We define the safety of parameters in terms of another a priori unknown and expensive to evaluate safety constraint $s\mathrel{\mathop{:}}\mathcal{X}\rightarrow\mathbb{R}$, s.t. parameters that satisfy $s(\bm{x})\geq 0$ are classified as safe, while the others as unsafe  [^1] This choice is without loss of generality, since we can incorporate any known non-zero safety threshold $\tau(\bm{x})$ in the constraint: $s^{\prime}(\bm{x})\coloneqq s(\bm{x})-\tau(\bm{x})$.. To start exploring safely, we also assume to have access to an initial safe parameter $\bm{x}_{0}$ that satisfies the safety condition, $s(\bm{x}_{0})\geq 0$.
@@
Figure 1: In (0(a)) we illustrate the safe optimization task. Based on the unknown safety constraint $s$, we are only allowed to evaluate parameters $\bm{x}$ with values $s(\bm{x})$ above the safety threshold (dashed line). Starting from a safe seed $\bm{x}_{0}$ a safe optimization strategy needs to find the optimum of the unknown objective $f$ within the largest reachable safe region of the parameter space containing $\bm{x}_{0}$. In (0(b)) we show the mutual information $I_{n}(\{\bm{x},y\};\mathbb{I}_{s(\cdot)\geq 0}(\bm{z}))$ in green for different parameters $\bm{x}$ inside the safe set and for a fixed $\bm{z}$ outside (red dashed line). The exploration part of our algorithm maximizes this quantity jointly over $\bm{x}$ and $\bm{z}$.
@@ 2.1 Probabilistic model
As both $f$ and $s$ are unknown and the evaluations $y_{n}^{f/s}$ are noisy, it is not feasible to select parameters that are safe with certainty. Instead, we provide high-probability safety guarantees by exploiting a probabilistic model for $f$ and $s$. To this end, we follow Berkenkamp et al. (2016a) and extend the domain $\mathcal{X}$ to $\tilde{\mathcal{X}}\coloneqq\mathcal{X}\times\{0,1\}$ and define the function $h\mathrel{\mathop{:}}\tilde{\mathcal{X}}\rightarrow\mathbb{R}$ as
@@
Furthermore, we assume that $h$ has bounded norm in the Reproducing Kernel Hilbert Spaces (RKHS) (Schölkopf and Smola, 2002) $\mathcal{H}_{k}$ associated to some kernel $k\mathrel{\mathop{:}}\tilde{\mathcal{X}}\times\tilde{\mathcal{X}}\rightarrow\mathbb{R}$ with $k\mathopen{}\mathclose{{}\left((\bm{x},i),(\bm{x}^{\prime},j)}\right)\leq 1$: $\mathinner{\!\mathopen{}\mathclose{{}\left\lVert h}\right\rVert}_{\mathcal{H}_{k}}=B<\infty$. This assumption allows us to use a Gaussian process (GP) (Srinivas et al., 2010) as probabilistic model for $h$.
A Gaussian process is a stochastic process specified by a mean function $\mu\mathrel{\mathop{:}}\tilde{\mathcal{X}}\rightarrow\mathbb{R}$ and a kernel $k$ (Rasmussen and Williams, 2006). It defines a probability distribution over real-valued functions on $\tilde{\mathcal{X}}$, such that any finite collection of function values at parameters $[(\bm{x}_{1},i_{1}),\dots,(\bm{x}_{n},i_{n})]$ is distributed as a multivariate normal distribution. The GP prior can then be conditioned on (noisy) function evaluations $\mathcal{D}_{n}=\{\mathopen{}\mathclose{{}\left((\bm{x}_{j},i_{j}),y_{j}}\right)\}_{j=1}^{n}$. If the noise at each observation is Gaussian, $\nu_{n}\sim\mathcal{N}(0,\sigma_{\nu}^{2})$ $\forall n$, then the resulting posterior is also a GP with posterior mean and variance given by
@@ 2.2 Safe set
$S_{n}\coloneqq\{\bm{x}\in\mathcal{X}\mathrel{\mathop{:}}\mu_{n}^{s}(\bm{x})-\beta_{n}\sigma_{n}^{s}(\bm{x})\geq 0\}\cup S_{n-1};~{}~{}S_{0}\doteq\bm{x}_{0},$ (6)
@@
Choose $\delta\in(0,1)$ and let $h$ as defined in 2 be an element in the RKHS $\mathcal{H}_{k}$ associated to some kernel $k$, with $\mathinner{\!\mathopen{}\mathclose{{}\left\lVert h}\right\rVert}_{\mathcal{H}_{k}}=B<\infty$. Let, moreover, the sequence of positive numbers $\{\beta_{n}\}$ be as in 5. Then if we define the safe set as in 6, we have that $s(\bm{x})\geq 0$ with probability of at least $1-\delta$, jointly for all $\bm{x}\in S_{n}$ and for all $n$.
@@ 2.4 Bayesian optimization
If the set $S_{\varepsilon}(\bm{x}_{0})$ were known, then the problem would be global optimization within $S_{\varepsilon}(\bm{x}_{0})$. A popular framework to find the global optimum of an unknown and expensive function via sequential evaluations is Bayesian optimization. Given a function $f$ defined on some fixed domain $\Omega$, BO algorithms find the $\operatorname*{arg\,max}_{\bm{x}\in\Omega}f(\bm{x})$ by building a probabilistic model of $f$ and evaluating parameters that maximize a so called acquisition function $\alpha\mathrel{\mathop{:}}\Omega\rightarrow\mathbb{R}$ defined via that model. The probabilistic model is then updated in a Bayesian fashion using the evaluation results. Algorithm 1 shows pseudo-code for the generic BO loop.
@@ 3.2 Optimization
Figure 3: Example of the two components of the acquisition function. For simplicity, here the objective $f$ and the constraint $s$ are the same function. The two plots show the values of the safe set expansion component ($\alpha^{\textsc{ISE}}$) and the optimization component ($\alpha^{\textsc{MES}}$) of the acquisition function inside the safe set ($S_{n}$), against the GP posterior mean and confidence interval (the blue curve and shaded area, respectively), with the safety threshold $s(\bm{x})=f(\bm{x})=0$ indicated by the orange dashed line. At early stages (2(a)) we see that both components achieve their maximum on the boundary of the safe set, since, according to the GP posterior, it is both promising as a region that contains the optimum and as a region that can give us much information about the safety of parameters outside $S_{n}$. The plot in (2(b)), however, shows that, as we continue sampling on the border, $\alpha^{\textsc{MES}}$ vanishes here, since it is unlikely that the optimum is in these neighborhoods. On the contrary, $\alpha^{\textsc{ISE}}$ remains non negligible on the boundaries, as these parameters can still give us information about the safety constraint outside of the current $S_{n}$, possibly leading to its expansion.
@@ 4.1 Definition of $S_{\varepsilon}(\bm{x}_{0})$
[…]
@@
[…]
@@ 4.2 Convergence results
As a first result, we show that by sampling only according to $\alpha^{\textsc{ISE}}$, i.e., $\bm{x}_{n+1}\in\operatorname*{arg\,max}_{\bm{x}\in S_{n}}\alpha^{\textsc{ISE}}(\bm{x})$, we eventually classify as safe the whole $S_{\varepsilon}(\bm{x}_{0})$. To this end, we introduce the set $\bar{S}(s)\coloneqq\mathopen{}\mathclose{{}\left\{\bm{x}\in\mathcal{X}\mathrel{\mathop{:}}s(\bm{x})\geq 0}\right\}$, which is the true safe set of the constraint $s$. With it, we can present the following result:
@@
Let the domain $\mathcal{X}$ be discrete and of size $D$: $|\mathcal{X}|=D<\infty$. Assume, moreover, that $\bm{x}_{n+1}$ is chosen according to $\bm{x}_{n+1}\in\operatorname*{arg\,max}_{\bm{x}\in S_{n}}\alpha^{\textsc{ISE}}(\bm{x})$. Then, if we define $N_{S}$ as the size of the true safe set of $s$, $N_{S}=|\bar{S}(s)|$, it holds that for all $\varepsilon>0$ there exists $N_{\varepsilon}$ such that, with probability of at least $1-\delta$, $S_{n}\supseteq S_{\varepsilon}(\bm{x}_{0})$ for all $n\geq N_{S}N_{\varepsilon}$. The smallest of such $N_{\varepsilon}$ is given by
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}\mathrel{\mathop{:}}\beta_{N}\eta^{-1}\mathopen{}\mathclose{{}\left(\frac{C\gamma_{N}}{N}}\right)\leq\varepsilon}\right\},$ (20)
@@
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}\mathrel{\mathop{:}}\beta_{N}b^{-1}\mathopen{}\mathclose{{}\left(\frac{C\gamma_{N}}{N}}\right)\leq\varepsilon}\right\},$ (21)
@@ 6 Experiments
Figure 5: Performance of ISE-BO compared to SafeOpt and both constrained and unconstrained versions of MES, when the objective and constraint function are samples drawn the a GP. In both (4(b)) and (4(a)) we can see the average simple regret 22 and its standard error for the tested methods over 50 random seeds. In (4(a)), the same GP sample served as both the objective $f$ and the constraint $s$, while in (4(b)), for each random seed, two different GP samples were drawn to serve, respectively, as the objective and the constraint.
@@
Figure 6: Performance of ISE-BO compared to SafeOpt and both constrained and unconstrained versions of MES for a one dimensional synthetic function, with $f=s$. In (5(a)), we plot this function, with the value of the safe seed indicated by the red cross and the one of the safe optimum by the pink square, while in (5(b)) we plot the average simple regret and its standard error over 50 random seeds. ISE-BO will first focus on the left of the safe seed. Once it has learned about the value of the optimum at $\bm{x}=-2.5$, it will it find more informative to explore to the right, eventually uncovering the region containing $f^{*}_{\varepsilon}$. On the other hand, MES-safe will spend most of its time around the local optimum at $\bm{x}=-2.5$. Similarly, SafeOpt algorithms with a too low or too high value of the Lipschitz constant will take longer to bypass the region where the constraint is very close to the safety threshold.
@@
Figure 7: Performance comparison in higher dimension with heteroskedastic noise, for $d=$4 and $d=$6, where we plot the simple regret $r_{n}$ with respect to the safe optimum. (6(a)) and (6(b)) show, respectively, the average regret over 30 runs in dimension four and six, as a function of the number of iterations. We can see that this adapted version of ISE-BO promotes expansion of the safe set, leading to classifying as safe regions where the latent function attains its largest value. The plots also show that ISE-BO achieves a better sample efficiency than both StageOpt-like exploration and the StageOpt inspired heuristic acquisition. As one would expect, we also see that the constrained version of MES achieves the worst performance, as it lacks motivation to expand the safe set. On the other hand, the unconstrained MES algorithm is naturally the quickest at identifying the optimum, since it is free to evaluate any parameter in the domain.
@@
After investigating the performance of ISE-BO on GP samples, we apply its exploration component to a classic control task from the OpenAI Gym framework (Brockman et al., 2016), with the goal of finding the set of parameters of a controller that satisfy some safety constraint. In particular, we consider a linear controller for the inverted pendulum. The controller is given by $u_{t}=x_{1}\theta_{t}+x_{2}\dot{\theta}_{t}$, where $u_{t}$ is the control signal at time $t$, while $\theta_{t}$ and $\dot{\theta}_{t}$ are, respectively, the angular position and the angular velocity of the pendulum. Starting from a position close to the upright equilibrium, the controller’s task is the stabilization of the pendulum, subject to a safety constraint on the maximum velocity reached within one episode. For a given initial controller configuration $\bm{x}_{0}\coloneqq(x_{1}^{0},x_{2}^{0})$, we want to explore the controller’s parameter space avoiding configurations that lead the pendulum to swing with a too high velocity. Namely, the safety constraint is the maximum angular velocity reached by the pendulum during an episode of fixed length $T$, $s(\bm{x})=\max_{t\leq T}\dot{\theta}_{t}(\bm{x})$, and the safety threshold is not at zero, but rather at some finite value $\dot{\theta}_{M}=0.5\text{rad/}s$, so that the safe parameters are those for which the maximum velocity is below $\dot{\theta}_{M}$. In Figure 7(a), we show the true safe set for this problem, while in Figures 7(b), 7(c) and 7(d) one can see how ISE safely explores the true safe set. These plots show the parameters that ISE chooses to evaluate at different iterations. We see that in all these occasions, the ISE acquisition function focuses on the boundary of the current safe set, as this is the region that is naturally most informative about the safety of parameters outside of the safe set. This way, ISE is able to expand the safe set, until it learns the constraints function with high confidence in the vicinity of the boundary, at which point the safe set can only change slightly. This behavior eventually leads to the full true safe set to be classified as safe by the GP posterior, as Figure 7(d) shows.
@@
Figure 8: Safe exploration of the linear controller’s parameter space in the inverted pendulum experiment. In (7(a)) we see the true safe set, while in (7(b)-7(d)) we see the safe set (blue region) as learned by ISE at various iterations. The point marked by the green dot is $\bm{x}_{n+1}=\operatorname*{arg\,max}\alpha^{\textsc{ISE}}$, while the yellow cross is the initial safe seed $\bm{x}_{0}$.

### v2

@@ 1.1 Related work
In the presence of constraints that the final solution needs to satisfy, but which we can violate during exploration, Gelbart et al. (2014) propose to combine typical BO acquisition functions with the probability of satisfying the constraint. Instead, Gotovos et al. (2013) propose an uncertainty-based criterion that learns about the feasible region of parameters, while Perrone et al. (2019) modify the Max-value Entropy Search selection criterion (Wang and Jegelka, 2017) to include information about the constraint as well in the acquisition function. When we are not allowed to ever evaluate unsafe parameters, safe exploration is a necessary sub-routine of BO algorithms to learn about the safety of parameters. To safely explore, Schreiter et al. (2015) globally learn about the constraint by evaluating the most uncertain parameters. SafeOpt by Sui et al. (2015) extends this to joint exploration and optimization and makes it more efficient by explicitly restricting safe exploration to the boundary of the safe set. Sui et al. (2018) proposes StageOpt, which additionally separates the exploration and optimization phases. Both of these algorithms assume access to a Lipschitz constant to define parameters close to the boundary of the safe set, which is a difficult tuning parameter in practice. These methods have been extended to multiple constraints by Berkenkamp et al. (2016a), while Kirschner et al. (2019) scale them to higher dimensions with LineBO, which explores in low-dimensional sub-spaces. To improve computational costs, Duivenvoorden et al. (2017) suggest a continuous approximation to SafeOpt without providing exploration guarantees. Chan et al. (2023), on the other hand, take a more directed approach, by searching the most promising parameter also outside of the current safe set and then evaluating the most uncertain safe parameter in the neighborhood of the projection of such promising parameter on the boundary of the safe set. In the context of safe active learning, Li et al. (2022) consider the safe active learning problem for a multi-output Gaussian process, and choose the parameters to evaluate as those with the highest predictive entropy and high probability of being safe. All of these methods rely on function uncertainty to drive exploration, while we directly maximize the information gained about the safety of parameters. In parallel independent work, Hübotter et al. (2024) proposed ITL, an active learning algorithm that aims at learning the values of a function in a target set $\mathcal{A}$ by actively sampling observations in a sample space $\mathcal{S}$, whose relation to $\mathcal{A}$ can be arbitrary. To do so, ITL maximizes the mutual information between the prediction targets in $\mathcal{A}$ and the next observation in $\mathcal{S}$. While addressing a more generic problem, when the target space is the set of potential maximizers and the sample space the set of safe parameters, ITL effectively becomes a safe BO algorithm.
@@ 2 Problem Statement and Preliminaries
In this section, we introduce the problem and the notation that we use throughout the paper. We are given an unknown and expensive function $f:\mathcal{X}\rightarrow\mathbb{R}$ that we aim to optimize by sequentially evaluating it at parameters $\bm{x}_{n}\in\mathcal{X}$. However, we are not allowed to evaluate parameters that violate some notion of safety. We define the safety of parameters in terms of another a priori unknown and expensive to evaluate safety constraint $s:\mathcal{X}\rightarrow\mathbb{R}$, s.t. parameters that satisfy $s(\bm{x})\geq 0$ are classified as safe, while the others as unsafe  [^1] This choice is without loss of generality, since we can incorporate any known non-zero safety threshold $\tau(\bm{x})$ in the constraint: $s^{\prime}(\bm{x})\coloneqq s(\bm{x})-\tau(\bm{x})$.. To start exploring safely, we also assume to have access to an initial safe parameter $\bm{x}_{0}$ that satisfies the safety condition, $s(\bm{x}_{0})\geq 0$.
@@
Figure 1: In (1(a)) we illustrate the safe optimization task. Based on the unknown safety constraint $s$, we are only allowed to evaluate parameters $\bm{x}$ with values $s(\bm{x})$ above the safety threshold (dashed line). Starting from a safe seed $\bm{x}_{0}$ a safe optimization strategy needs to find the optimum of the unknown objective $f$ within the largest reachable safe region of the parameter space containing $\bm{x}_{0}$. In (1(b)) we show the mutual information $I_{n}(\{\bm{x},y\};\mathbb{I}_{s(\cdot)\geq 0}(\bm{z}))$ in green for different parameters $\bm{x}$ inside the safe set and for a fixed $\bm{z}$ outside (red dashed line). The exploration part of our algorithm maximizes this quantity jointly over $\bm{x}$ and $\bm{z}$.
@@ 2.1 Probabilistic model
As both $f$ and $s$ are unknown and the evaluations $y_{n}^{f/s}$ are noisy, it is not feasible to select parameters that are safe with certainty. Instead, we provide high-probability safety guarantees by exploiting a probabilistic model for $f$ and $s$. To this end, we follow Berkenkamp et al. (2016a) and extend the domain $\mathcal{X}$ to $\tilde{\mathcal{X}}\coloneqq\mathcal{X}\times\{0,1\}$ and define the function $h:\tilde{\mathcal{X}}\rightarrow\mathbb{R}$ as
@@
[…]
A Gaussian process is a stochastic process specified by a mean function $\mu:\tilde{\mathcal{X}}\rightarrow\mathbb{R}$ and a kernel $k$ (Rasmussen and Williams, 2006). It defines a probability distribution over real-valued functions on $\tilde{\mathcal{X}}$, such that any finite collection of function values at parameters $[(\bm{x}_{1},i_{1}),\dots,(\bm{x}_{n},i_{n})]$ is distributed as a multivariate normal distribution. The GP prior can then be conditioned on (noisy) function evaluations $\mathcal{D}_{n}=\{\mathopen{}\mathclose{{}\left((\bm{x}_{j},i_{j}),y_{j}}\right)\}_{j=1}^{n}$. If the noise at each observation is Gaussian, $\nu_{n}\sim\mathcal{N}(0,\sigma_{\nu}^{2})$ $\forall n$, then the resulting posterior is also a GP with posterior mean and variance given by
@@ 2.2 Safe set
$S_{n}\coloneqq\{\bm{x}\in\mathcal{X}:\mu_{n}^{s}(\bm{x})-\beta_{n}\sigma_{n}^{s}(\bm{x})\geq 0\}\cup S_{n-1};~{}~{}S_{0}\doteq\bm{x}_{0},$ (6)
@@
[…]
@@ 2.4 Bayesian optimization
If the set $S_{\varepsilon}(\bm{x}_{0})$ were known, then the problem would be global optimization within $S_{\varepsilon}(\bm{x}_{0})$. A popular framework to find the global optimum of an unknown and expensive function via sequential evaluations is Bayesian optimization. Given a function $f$ defined on some fixed domain $\Omega$, BO algorithms find the $\operatorname*{arg\,max}_{\bm{x}\in\Omega}f(\bm{x})$ by building a probabilistic model of $f$ and evaluating parameters that maximize a so called acquisition function $\alpha:\Omega\rightarrow\mathbb{R}$ defined via that model. The probabilistic model is then updated in a Bayesian fashion using the evaluation results. Algorithm 1 shows pseudo-code for the generic BO loop.
@@ 3.2 Optimization
Figure 3: Example of the two components of the acquisition function. For simplicity, here the objective $f$ and the constraint $s$ are the same function. The two plots show the values of the safe set expansion component ($\alpha^{\textsc{ISE}}$) and the optimization component ($\alpha^{\textsc{MES}}$) of the acquisition function inside the safe set ($S_{n}$), against the GP posterior mean and confidence interval (the blue curve and shaded area, respectively), with the safety threshold $s(\bm{x})=f(\bm{x})=0$ indicated by the orange dashed line. At early stages (3(a)) we see that both components achieve their maximum on the boundary of the safe set, since, according to the GP posterior, it is both promising as a region that contains the optimum and as a region that can give us much information about the safety of parameters outside $S_{n}$. The plot in (3(b)), however, shows that, as we continue sampling on the border, $\alpha^{\textsc{MES}}$ vanishes here, since it is unlikely that the optimum is in these neighborhoods. On the contrary, $\alpha^{\textsc{ISE}}$ remains non negligible on the boundaries, as these parameters can still give us information about the safety constraint outside of the current $S_{n}$, possibly leading to its expansion.
@@ 4.1 Definition of $S_{\varepsilon}(\bm{x}_{0})$
As we explained in Section 2.3, the set $S_{\varepsilon}(\bm{x}_{0})$ is the set of all parameters that are safe with high probability and that are safely reachable from $\bm{x}_{0}$ by all well behaved GPs. In order to formally explain what we mean by it, we first need some preliminary definitions. In the following, we refer to a GP $\sim\mathcal{N}(\mu_{0},k)$ conditioned on some observation data $\mathcal{D}=\{(\bm{x}_{n},y_{n})\}$ as GPD. Using this notation, we can define the set of all well behaved GPs that have small uncertainty over a given set $\Omega$:
@@
$\begin{split}\bm{\beta}\text{-GP}_{s}^{\varepsilon}(\Omega)=\big{\{}\text{GP}_{\mathcal{D}}:&~{}s(\bm{x})\in[\mu_{n}(\bm{x})-\beta_{n}\sigma_{n}(\bm{x}),\mu_{n}(\bm{x})+\beta_{n}\sigma_{n}(\bm{x})]\forall\bm{x}\in\mathcal{X},\forall n\leq|\mathcal{D}|\big{.}\\ \big{.}&\wedge\beta_{|\mathcal{D}|}\sigma_{|\mathcal{D}|}(\bm{x})\leq\varepsilon~{}\forall\bm{x}\in\Omega\big{\}}.\end{split}$ (17)
As a consequence of the choice of the sequence $\mathopen{}\mathclose{{}\left\{\beta_{n}}\right\}$, we know that every GPD is well behaved with probability at least $1-\delta$, so that Definition 2 is not trivial (see Lemma 4 in Appendix A for more details).
@@
Given a dataset $\mathcal{D}$ and a GP conditioned on it, GPD, we call $S_{\text{GP}_{\mathcal{D}}}$ the safe set associated to such GP posterior, as prescribed by 6. With this notation, given a safe set $S$ and $\varepsilon>0$, we define the expansion operator $R_{\varepsilon}$ as
$R_{\varepsilon}(S)=\big{\{}\bm{x}\in\mathcal{X}:\bm{x}\in S_{\text{GP}_{\mathcal{D}}}\text{ for all GP }\in\bm{\beta}\text{-GP}_{s}^{\varepsilon}(S)\big{\}},$ (18)
@@ 4.2 Convergence results
As a first result, we show that by sampling only according to $\alpha^{\textsc{ISE}}$, i.e., $\bm{x}_{n+1}\in\operatorname*{arg\,max}_{\bm{x}\in S_{n}}\alpha^{\textsc{ISE}}(\bm{x})$, we eventually classify as safe the whole $S_{\varepsilon}(\bm{x}_{0})$. To this end, we introduce the set $\bar{S}(s)\coloneqq\mathopen{}\mathclose{{}\left\{\bm{x}\in\mathcal{X}:s(\bm{x})\geq 0}\right\}$, which is the true safe set of the constraint $s$. With it, we can present the following result:
@@
Let the domain $\mathcal{X}$ be discrete and of size $D$: $|\mathcal{X}|=D<\infty$. Assume, moreover, that $\bm{x}_{n+1}$ is chosen according to $\bm{x}_{n+1}\in\operatorname*{arg\,max}_{\bm{x}\in S_{n}}\alpha^{\textsc{ISE}}(\bm{x})$. Then, if we define $N_{S}$ as the size of the true safe set of $s$, $N_{S}=|\bar{S}(s)|$, it holds that for all $\varepsilon>0$ there exists $N_{\varepsilon}$ such that, with probability of at least $1-\delta$, $S_{n}\supseteq S_{\varepsilon}(\bm{x}_{0})$ for all $n\geq N_{S}N_{\varepsilon}$.
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}:\beta_{N}\eta^{-1}\mathopen{}\mathclose{{}\left(\frac{C\gamma_{N}}{N}}\right)\leq\varepsilon}\right\},$ (20)
@@
$N_{\varepsilon}=\min\mathopen{}\mathclose{{}\left\{N\in\mathbb{N}:\beta_{N}b^{-1}\mathopen{}\mathclose{{}\left(\frac{C\gamma_{N}}{N}}\right)\leq\varepsilon}\right\},$ (21)
@@ 6 Experiments
Figure 5: Performance of ISE-BO compared to SafeOpt and both constrained and unconstrained versions of MES, when the objective and constraint function are samples drawn the a GP. In both (5(b)) and (5(a)) we can see the average simple regret 22 and its standard error for the tested methods over 50 random seeds. In (5(a)), the same GP sample served as both the objective $f$ and the constraint $s$, while in (5(b)), for each random seed, two different GP samples were drawn to serve, respectively, as the objective and the constraint.
@@
Figure 6: Performance of ISE-BO compared to SafeOpt and both constrained and unconstrained versions of MES for a one dimensional synthetic function, with $f=s$. In (6(a)), we plot this function, with the value of the safe seed indicated by the red cross and the one of the safe optimum by the pink square, while in (6(b)) we plot the average simple regret and its standard error over 50 random seeds. ISE-BO will first focus on the left of the safe seed. Once it has learned about the value of the optimum at $\bm{x}=-2.5$, it will it find more informative to explore to the right, eventually uncovering the region containing $f^{*}_{\varepsilon}$. On the other hand, MES-safe will spend most of its time around the local optimum at $\bm{x}=-2.5$. Similarly, SafeOpt algorithms with a too low or too high value of the Lipschitz constant will take longer to bypass the region where the constraint is very close to the safety threshold.
@@
Figure 7: Performance comparison in higher dimension with heteroskedastic noise, for $d=$4 and $d=$6, where we plot the simple regret $r_{n}$ with respect to the safe optimum. (7(a)) and (7(b)) show, respectively, the average regret over 30 runs in dimension four and six, as a function of the number of iterations. We can see that this adapted version of ISE-BO promotes expansion of the safe set, leading to classifying as safe regions where the latent function attains its largest value. The plots also show that ISE-BO achieves a better sample efficiency than both StageOpt-like exploration and the StageOpt inspired heuristic acquisition. As one would expect, we also see that the constrained version of MES achieves the worst performance, as it lacks motivation to expand the safe set. On the other hand, the unconstrained MES algorithm is naturally the quickest at identifying the optimum, since it is free to evaluate any parameter in the domain.
@@
After investigating the performance of ISE-BO on GP samples, we apply its exploration component to a classic control task from the OpenAI Gym framework (Brockman et al., 2016), with the goal of finding the set of parameters of a controller that satisfy some safety constraint. In particular, we consider a linear controller for the inverted pendulum. The controller is given by $u_{t}=x_{1}\theta_{t}+x_{2}\dot{\theta}_{t}$, where $u_{t}$ is the control signal at time $t$, while $\theta_{t}$ and $\dot{\theta}_{t}$ are, respectively, the angular position and the angular velocity of the pendulum. Starting from a position close to the upright equilibrium, the controller’s task is the stabilization of the pendulum, subject to a safety constraint on the maximum velocity reached within one episode. For a given initial controller configuration $\bm{x}_{0}\coloneqq(x_{1}^{0},x_{2}^{0})$, we want to explore the controller’s parameter space avoiding configurations that lead the pendulum to swing with a too high velocity. Namely, the safety constraint is the maximum angular velocity reached by the pendulum during an episode of fixed length $T$, $s(\bm{x})=\max_{t\leq T}\dot{\theta}_{t}(\bm{x})$, and the safety threshold is not at zero, but rather at some finite value $\dot{\theta}_{M}=0.5\text{rad/}s$, so that the safe parameters are those for which the maximum velocity is below $\dot{\theta}_{M}$. In Figure 8(a), we show the true safe set for this problem, while in Figures 8(b), 8(c) and 8(d) one can see how ISE safely explores the true safe set. These plots show the parameters that ISE chooses to evaluate at different iterations. We see that in all these occasions, the ISE acquisition function focuses on the boundary of the current safe set, as this is the region that is naturally most informative about the safety of parameters outside of the safe set. This way, ISE is able to expand the safe set, until it learns the constraints function with high confidence in the vicinity of the boundary, at which point the safe set can only change slightly. This behavior eventually leads to the full true safe set to be classified as safe by the GP posterior, as Figure 8(d) shows.
@@
Figure 8: Safe exploration of the linear controller’s parameter space in the inverted pendulum experiment. In (8(a)) we see the true safe set, while in (8(b)-8(d)) we see the safe set (blue region) as learned by ISE at various iterations. The point marked by the green dot is $\bm{x}_{n+1}=\operatorname*{arg\,max}\alpha^{\textsc{ISE}}$, while the yellow cross is the initial safe seed $\bm{x}_{0}$.